# Field-free, Quasi-continuous Operation of Optical Nanofiber Interface with Two-dimensional Ferromagnetic Trap

### v3

@@ IV Toward an ultra-straight 2D ferromagnetic trap
As in Fig. 2(a,b), the field zero in the two-foil assembly is a 3D one. The field variation along $z$ only vanishes in the $\varepsilon=h/L\ll 1$ limit. On the other hand, this field variation is absent in the four-foil version, as being described in Fig. 2(c,d). Of course, the assumption is that the ${\bf\mu}-$metal foils are perfectly magnetized and aligned. In this section we check practical aspects of this assumption, by measuring the field distribution and compare with numerical simulations.
@@ IV.1 Field uniformity and stability
The field measurement is performed with a Gauss probe with 1 mm sensor size and 0.50.50.5~{}0.5G resolution. The Gauss probe is mounted on a 3D translation stage, centered to the $S_{2}$ surface of the foil to be measured, and is gently pressed against the surface to fix the distance. Then, as in Fig. 5a, by translating the Gauss probe along $z$, the surface field is recorded across the $L=120~{}$mm foil length. The measurements are repeated over days to verify the field stability. Example data are given in Fig. 5b with diamond, triangular, and disk markers. The marker size is chosen according to the measurement uncertainty and our confidence in positioning the Gauss probe.
@@
The measurement results are repeatable. Taking the 1-Foil-C data as an example, which consists of three repeated measurements conducted over the span of a week, the data deviation is less than 1 G, showing excellent overlap in the plot. On the other hand, given that NdFeB blocks exhibit variations in surface quality and magnetic strength, the overall surface field strength naturally varies from foil to foil, as illustrated in the examples in Fig. 5b. These variations can be managed by hand-picking NdFeB blocks with consistent properties or by fine-tuning the field strength using additional magnets.
We verify that the uniform field profile shown in Fig. 5b is insensitive to the way the source magnets are attached on the $S_{1}$ surface. Field variations at the $1\%$ level only occur near $S_{2}$ when one of the two NdFeB blocks is displaced on the $S_{1}$ surface by as much as a few millimeters—a displacement much larger than the positioning precision achievable with a millimeter-scale ruler. As discussed in Sec. II.1, the combination of efficient source-field shielding and surface field energy minimization enforces a uniform field distribution near $S_{2}$. For foils with large enough $W$, $L$ and $d$, field irregularities induced by variations of small magnets on the distant surface are suppressed. Similar shielding effects are expected to smooth out the irregularity in the foil-coil current [19].
Given the highly uniform surface field in the finely machined 4-foil assembly, the field distribution above $S_{2}$ for $h,h^{\prime}\ll L$ are also expected to be uniform (Fig. 2(c,d)), as dictated by field continuity [31]. In our prototype assembly, accurately mapping the 3D field distribution is technically challenging. Nevertheless, using the same Gauss probe, we measured the field both above the foil surfaces and near the zero-field region of the 4-foil assembly. These measurements align with simulations, albeit with reduced fractional resolution and increased positional uncertainties.
Finally, as in Sec. II.1, we emphasize the choice of large enough working distances $h$ for finely machined ${\bf\mu}-$metal foils, to avoid fringe fields associated with micrometer-scale ${\bf\mu}-$metal surface corrugations [27, 28].
@@ IV.2 Alignment considerations
The construction of the 2D quadrupole field shown in Fig. 2(c,d) relies on precise alignment of the four uniformly magnetized ${\bf\mu}$-metal foils. In practice, field deviations can arise from misalignment or positional errors. However, because the ${\bf\mu}$-metal foils can be accurately machined and securely mounted, we anticipate small alignment or positional errors in the foil assembly will be manageable to induce negligible field deviations.
[…]
@@
Next, we consider the effect of a positional error $\delta z$ along the $z$-axis by one of the foils. The resulting field perturbation can be interpreted as a shift in surface magnetic charge [25]. The unbalanced magnetic charge at the edges lead to a residual field at the center, which, for sufficiently large foil length $L$, becomes weak and smooth. Therefore, for precisely positioned foils with $\delta z/L$ at e.g. $10^{-4}$ level, and in addition $L\gg h$, this type of perturbation can be suppressed to milli-Gauss level while maintaining the $B_{1}\sim 100~{}$G/cm gradient.
@@
On the other hand, for any one of the foils, small positional errors along the $x$ and $y$ directions shifts the zero-field line itself, and can be compensated by fine-tuning the magnetization through adjustments to the foil-coil current.
IV.3 Field-free ONF interface with $l\sim 100~{}$mm
@@
So far, we have shown that it is practical to construct a 2D ferromagnetic trap as those in Fig. 2(c,d) with high precision. Enhanced by permanent magnets, a field gradient of $B_{1}=100~{}$G/cm can be quite easily achieved at $h=20~{}$mm distance for a four-foil assembly with e.g. $L=200$ mm. The center zero-field line is long enough to support ONF-2D-MOT system with $l\sim 100~{}$mm, with residual field suppressed to the milli-Gauss level. As such, the evanescent optical lattice [10, 11, 12, 13, 14] can be loaded from the high-gradient 2D-MOT without switching off the trapping field.
@@
For a uniformly filled lattice, the collective interaction strength between guided photons with the lattice atoms increases linearly with $l$. With the $l\approx 100~{}$mm interaction length, the interaction strength would be quite unprecedented, which can be crucial for uncovering new regime in waveguide QED [43, 2]. Importantly, the quantum optical measurements can be performed in a quasi-continuous manner, as those demonstrated in this work (Fig. 4), leading to high data rate. The efficient laser-cooling in between measurements should also help to mitigate ONF vibration induced heating [48].
We finally note that while the magnetization and alignment can be optimized to improve the field-free operation of the ONF-interface, practically the level of field suppression along ONF is still likely limited by engineering imperfections. To this end, we highlight previous successful experience on ferromagnetic trap with guided matterwave interferometry [19]. We also note that for achieving improved interaction uniformity, the inhomogeneous broadening due to weak residual fields can be managed by spin echo and composite pulse techniques [49, 37].
V Summary and outlook
Integrating nanophotonics with cold atoms represents a promising route toward quantum information processing via nonlinear quantum optics [50]. A major technical challenge in this endeavor is the precise positioning of cold atoms at the nanophotonic interface. Among the most successful approaches to address this challenge is the ONF-cold atom interface, which has seen significant progress over the last decade [16, 10, 8, 2, 3, 17, 6]. In these works, sub-Doppler cooling is critical for efficient loading of the near-field lattice, while a uniform atomic spectroscopic response is essential to support long-range resonant dipole interactions. But the associated field-free requirement is at odds with the necessity of a strong field gradient for magneto-optical trapping. Managing these conflicting requirements becomes increasingly difficult at large $l$, the ONF interaction distance, which is preferred to be as long as possible. But switching off the field to accommodate these constraints severely limits the duty cycle of single-photon-level measurements.
In this work, we have proposed to resolve the contradictory field requirement at the cold-atom-ONF platform using 2D ferromagnetic trap [18, 19]. We highlight the unique opportunity of aligning the zero-field line in 2D-MOT with the stiff ONF, for efficient loading and field-free operation of the nanophotonic interface. We show that precisely fabricated ${\bf\mu}-$metal foils robustly uniformize source field from permanent magnets or current-carrying wires, for generating ultra-straight 2D field to match a long and straight ONF [5].
Experimentally, we integrate a two-foil ferromagnetic assembly with ONF using a planar 2D-MOT (Fig. 3). The field uniformity is confirmed with Gauss probe measurement on the foil surface (Fig. 5b). The 2D nature of the central quadrupole field (Fig. 2(a,b)) is verified by the robust operation of the highly elongated 2D MOT (Fig. 3e). Freed from the slow field-switching constraints, we perform atomic spectroscopy at a repetition rate as high as $f_{\rm rep}=250~{}$kHz, utilizing nanosecond probe pulses as shown in Fig. 4. We note that the quasi-continuous spectroscopic measurements minimally perturb the atomic trajectories [10]. In the ONF near field, the atomic density distribution is preserved as that during MOT operation, a situation not been carefully looked upon with precision spectroscopy previously, to our best knowledge. Therefore, the anomalous line-broadening effect in this unique system, observed in Sec. III.5, warrants further investigations [47].
Comparing with traditional ONF-interfaces, the ferromagnetic, planar 2D-MOT interface demonstrated in this work (Fig. 3a) is substantially simpler and more compact. The $\sim$20 G/cm gradient with sufficient tunability (Fig. 3f) is achieved at a notable long working distance of $h=35~{}$mm, facilitating the ex-vacuo operation. Based on field measurements and simulations, we provide arguments in Sec. IV that ONF interface with $l\approx 100~{}$mm can be constructed using a four-foil ferromagnetic trap. The unprecedentedly high light-atom coupling strength may help to unlock new research opportunities in waveguide QED [43, 2]. The quantum optical measurements can be performed in a quasi-continuous manner, as those demonstrated in this work. The efficient laser-cooling in between measurements should also help to mitigate ONF vibration induced heating [48].
@@ V Summary and outlook
Separately, with reduced $h$, the field gradient in these ferromagnetic traps can easily reach $B_{1}=100$ G/cm, strong enough for tightly confining laser-cooled atoms in their low-field-seeking states with magnetic dipole forces [19]. This may open up new quantum optical scenarios featuring waveguided photons interacting with magnetically guided atoms. To this end, we note related progresses are attempted with optical guiding inside hollow-core fibers [51, 52], where the field inhomogenuity poses the key challenge. Here, the ultra-smooth confining potential offered by the ferromagnetic trap would help to support matterwave coherence over a long distance comparable to that for the optical coherence in the ONF-based quantum optical platform [2, 3].
VI Acknowledgments
@@ Appendix A Magnetic field simulation
The accuracy of the field simulation relies on the granularity of magnet segmentation and the numerical relaxation tolerance. We adjusted these parameters to ensure that neither finer subdivisions nor tighter relaxation tolerances significantly affected the final results. The relaxation tolerance is set to ${\rm prec}=10^{-4}$ T (1 G). For the simulations in Fig. 2 and Fig. 5c, each NdFeB block is divided into $10\times 10\times 10$ subunits. Each ${\rm{\bf\mu}}$-metal foil is divided into $n_{z}=160$ and $n_{y}=50$ segments along the $z$ (“$L$ direction”) and $y$ (“$W$ direction”) axes, respectively. Additional simulations confirmed that the central field profile $B_{y}(z)$, as shown in Fig. 5c, is insensitive to segmentation along the $x$-axis, allowing us to set $n_{x}=1$ to conserve computational resources. This approach neglects the variation of $B_{y}$ across the foil thickness ($x$), which is minimized at the center ($x=0$) and peaks at the edges. To account for this approximation, an overall scaling factor of 0.8 is applied to the field simulation in Fig. 5c, evaluated at $h=0.5~{}\mathrm{mm}$ from $S_{2}$, to match experimental measurements shown in Fig. 5b. This correction is necessary only in the vicinity of $S_{2}$. For field distributions where $h\gg d$, the simulation remains accurate, and we fix the NdFeB block surface field at $B_{0}=0.9~{}\mathrm{T}$ for all simulations.

### v5

@@ IV Toward an ultra-straight 2D ferromagnetic trap
As shown in Fig. 2(a,b), the field zero in the two-plate assembly is inherently three-dimensional. The field variation along $z$ only vanishes in the limit $\varepsilon=h/L\ll 1$. On the other hand, this field variation is absent in the four-plate configuration, as described in Fig. 2(c,d). Of course, this ideal behavior assumes that the $\bf\mu$}-metal plates are perfectly magnetized and precisely aligned. In this section, we examine practical aspects of this assumption by measuring the field distribution and comparing it with numerical simulations.
Before proceeding with the discussion, it is worth noting that the highly uniform field distribution shown in Fig. 2(c,d) arises from the symmetry of the source field, which is robustly enforced by the magnetic response (Sec. II.1) of the precisely aligned $\mu$-metal plates. In principle, comparable uniformity could be achieved using current-carrying wires alone [51]. In practice, however, achieving this is technically challenging: aside from the difficulty of perfect winding, the current flow within conductors does not always conform precisely to the wire geometry [27, 52].
@@ IV.1 Field uniformity and stability
The field measurement is performed with a Gauss probe with 1 mm sensor size and 0.50.5~G resolution. The Gauss probe is mounted on a 3D translation stage, centered to the $S_{2}$ surface of the plate to be measured, and is gently pressed against the surface to fix the distance. Then, as in Fig. 5a, by translating the Gauss probe along $z$, the surface field is recorded across the L=120L=120~$m$ plate length. The measurements are repeated over days to verify the field stability. Example data are given in Fig. 5b with diamond, triangular, and disk markers. The marker size is chosen according to the measurement uncertainty and our confidence in positioning the Gauss probe.
@@
The measurement results are repeatable. Taking the 1-plate-C data as an example, which consists of three repeated measurements conducted over the span of a week, the data deviation is less than 1 G, showing excellent overlap in the plot. On the other hand, given that NdFeB blocks exhibit variations in surface quality and magnetic strength, the overall surface field strength naturally varies from plate to plate, as illustrated in the examples in Fig. 5b. These variations can be managed by hand-picking NdFeB blocks with consistent properties or by fine-tuning the field strength using additional magnets.
We verify that the uniform field profile shown in Fig. 5b is insensitive to the way the source magnets are attached on the $S_{1}$ surface. Field variations at the $1\%$ level only occur near $S_{2}$ when one of the two NdFeB blocks is displaced on the $S_{1}$ surface by as much as a few millimeters—a displacement much larger than the positioning precision achievable with a millimeter-scale ruler. As discussed in Sec. II.1, the combination of efficient source-field shielding and surface field energy minimization enforces a uniform field distribution near $S_{2}$. For plates with large enough $W$, $L$ and $d$, field irregularities induced by variations of small magnets on the distant surface are suppressed. Similar shielding effects are expected to smooth out the irregularity in the plate-coil current [19].
Given the highly uniform surface field in the finely machined 4-plate assembly, the field distribution above $S_{2}$ for $h,h^{\prime}\ll L$ are also expected to be uniform (Fig. 2(c,d)), as dictated by field continuity [31]. In our prototype assembly, accurately mapping the 3D field distribution is technically challenging. Nevertheless, using the same Gauss probe, we measured the field both above the plate surfaces and near the zero-field region of the 4-plate assembly. These measurements align with simulations, albeit with reduced fractional resolution and increased positional uncertainties.
As in Sec. II.1, we emphasize the choice of large enough working distances $h$ for finely machined $\bf\mu$}-metal plates, to avoid fringe fields associated with micrometer-scale $\bf\mu$}-metal surface corrugations [27, 28].
@@ IV.2 Alignment considerations
The construction of the 2D quadrupole field shown in Fig. 2(c,d) relies on precise alignment of the four uniformly magnetized $\bf\mu$}-metal plates. In practice, field deviations can arise from misalignment or positional errors. However, because the $\bf\mu$}-metal plates can be accurately machined and securely mounted, we anticipate small alignment or positional errors in the plate assembly will be manageable to induce negligible field deviations.
First, when the four plates are not perfectly parallel, then, depending on the kind of relative angular misalignment, the zero-field line is affected in different ways. However, with the rigidity provided by the precisely machined $\bf\mu$}-metal plates, we estimate that angular alignment precision can reach a 10-micro-radian level. For 2D trap with moderate field gradients, e.g. $B_{1}\sim 100~\text{G/cm}$, the residual field at center within a l=l=~100 mm distance due to the angular errors would be limited to 10 milli-Gauss level, regardless of relative orientations among the plates.
@@
Next, we consider the effect of a positional error $\delta z$ along the $z$-axis by one of the plates. The resulting field perturbation can be interpreted as a shift in surface magnetic charge [25]. The unbalanced magnetic charge at the edges leads to a residual field at the center, which, for sufficiently large plate length $L$, becomes weak and smooth. Therefore, for precisely positioned plates with $\delta z/L$ at e.g. $10^{-4}$ level, and in addition $L\gg h$, this type of perturbation can be suppressed to milli-Gauss level while maintaining the $B_{1}\sim 100$~G/cm gradient.
@@
On the other hand, for any one of the plates, small positional errors along the $x$ and $y$ directions shifts the zero-field line itself, and can be compensated by fine-tuning the magnetization through adjustments to the plate-coil current.
IV.3 Field-free ONF interface with $l\sim 100$~$m$
@@
So far, we have shown that it is practical to construct a 2D ferromagnetic trap as those in Fig. 2(c,d) with high precision. Enhanced by permanent magnets, a field gradient of $B_{1}=100$~G/cm can be quite easily achieved at h=20h=20~$m$ distance for a four-plate assembly with e.g. $L=200$ mm. The center zero-field line is long enough to support ONF-2D-MOT system with $l\sim 100$~$m$, with residual field suppressed to the milli-Gauss level. As such, the evanescent optical lattice [10, 11, 12, 42, 14] can be loaded from the high-gradient 2D-MOT without switching off the trapping field.
@@
For a uniformly filled lattice, the collective interaction strength between guided photons with the lattice atoms increases linearly with $l$. Scaling up from an existing apparatus [42], with the $l\approx 100$~$m$ interaction length, the resonant optical depth for an atomic array would reaches ${\rm OD}\sim 10^{3}$ level. This interaction strength would be quite unprecedented, which can be crucial for uncovering new regime in waveguide QED [45, 2]. Importantly, the quantum optical measurements can be performed in a quasi-continuous manner, as those demonstrated in this work (Fig. 4), leading to high data rate. The efficient laser-cooling in between measurements should also help to mitigate ONF vibration induced heating [53].
V Discussions
V.1 Magnetic guiding
So far, with nanosecond ONF transmission spectroscopy for the $n=2$ assembly (Sec. III.4) and a field-characterization study for the $n=4$ version (Sec. IV), our discussions have been focused on how the ultra-smooth field generated by the ferromagnetic assembly improves the ONF interface technology for free [17, 8] and lattice-confined atoms [10, 11, 12, 42, 14].
@@ V.1 Magnetic guiding
Clearly, with reduced $h$, the field gradient in these ferromagnetic traps can easily reach $B_{1}=100$ G/cm or higher, strong enough for tightly confining laser-cooled atoms in their low-field-seeking states with magnetic dipole forces [19]. Spin flips near the zero-field line may be suppressed by the vector shift of ONF guided light in the near field. Alternatively, taking advantage of the fast $\mu$-metal response [26], one may create a two-dimensional time-orbiting potential trap [54], by periodically driving the offset currents at high enough frequencies e.g. near 100 kHz. The setup may open up new quantum optical scenarios featuring waveguided photons interacting with magnetically guided atoms. To this end, we note related progresses are attempted with optical guiding inside hollow-core fibers [55, 56], where the field inhomogenuity poses the key challenge. Here, the ultra-smooth confining potential [19] offered by the ferromagnetic trap would help to support matterwave coherence over a long distance comparable to that for the optical coherence in the ONF-based quantum optical platform [2, 3].
V.2 Beyond $n=4$: Multipole fields
The soft ferromagnetic assembly can be conveniently constructed to support various atomic cooling and trapping geometries [57, 19]. In particular, the linear quadruple field in Fig. 2 can be extended to higher-order trapping fields, such as the $n=6$ hexapole (Fig. 6a) and $n=8$ octupole (Fig. 6b) configurations. Similar to the $n=4$ configuration, the higher-order fields are generated by $\nu=n/2\geq 3$ pairs of polarization-balanced magnets. The field variation along $z$ can therefore be efficiently suppressed to support highly uniform zero-field lines at the assembly center, as in Sec. IV. These high-order fields can be quite useful to extend the functionality of the ONF interface. For example, since the field strength around the higher-order field zero scales as $|B(\rho)|\propto\rho^{\nu-1}$ for $\rho=\sqrt{x^{2}+y^{2}}$, a higher-order trap provides a larger field-free volume [58] around ONF during laser cooling.
Furthermore, as schematically illustrated in Fig. 6, by adjusting the relative strengths of ferromagnetic plate pairs, the high-order zero-field line can be split into multiple low-order zero-field lines. Taking the hexapole field in Fig. 6a as an example. Here, the $P_{5,6}$-plates must be poled with sufficient strength to counter the $B_{x}$ field at the assembly center generated by the $P_{1,2,3,4}$ plates. A hexapole zero is formed only at a critical $P_{5,6}$ poling strength, beyond which the 2nd-order zero is split into two 1st-order zeroes to the left and right. Similarly, for the octupole field in Fig. 6b, by ramping up the poling strength of $P_{5,6,7,8}$ plates, the 3rd-order zero at the center is split into three first-order zeroes. The ability to split and merge multiple zero-field lines, combined with the ultra-smooth magnetic guiding (Sec. V.1), may enable coherent control of multiple 1D samples around ONF to interact with the guided photons.
Figure 6: (a) Schematic of an $n=6$ assembly. The central open circle represents the hexapole field when the magnetization of the $P_{5,6}$ plates are set at a critical value, beyond which the second-order field zero is split into two two first-order zeroes, moving toward $P_{1,2}$ and $P_{3,4}$, as represented by the blue disks. (b) Schematic of an $n=8$ assembly which creates an octupole field when $P_{1}$ to $P_{8}$ share the same strength. By ramping up $P_{5\sim 8}$ currents to increase their magnetization, the 3rd-order zero at center is split into three 1st-order zeroes. The splitting and merging of the multiple zero-field lines enables coherent control of multiple 1D atomic samples around ONF.
V.3 Summary and outlook
Integrating nanophotonics with cold atoms represents a promising route toward quantum information processing via nonlinear quantum optics [59, 4]. A major technical challenge in this endeavor is the precise positioning of cold atoms at the nanophotonic interface. Among the most successful approaches to address this challenge is the ONF-cold atom interface, which has seen significant progress over the last decade [10, 16, 8, 2, 3, 17, 60, 33]. In these works, sub-Doppler cooling is critical for efficient loading of the near-field lattice, while a uniform atomic spectroscopic response is essential to support long-range resonant dipole interactions. But the associated field-free requirement is at odds with the necessity of a strong field gradient for magneto-optical trapping. Managing these conflicting requirements becomes increasingly difficult at large $l$, the ONF interaction distance, which is preferred to be as long as possible. But switching off the field to accommodate these constraints severely limits the duty cycle of single-photon-level measurements.
In this work, we have proposed to resolve the contradictory field requirement at the cold-atom-ONF platform using 2D ferromagnetic trap [18, 19]. We highlight the unique opportunity of aligning the zero-field line in 2D-MOT with the stiff ONF, for efficient loading and field-free operation of the nanophotonic interface. We show that precisely fabricated $\bf\mu$}-metal plates robustly uniformize source field from permanent magnets or current-carrying wires, for generating ultra-straight 2D field to match a long and straight ONF [5].
Experimentally, we integrate a two-plate ferromagnetic assembly with ONF using a planar 2D-MOT (Fig. 3). The field uniformity is confirmed with Gauss probe measurement on the plate surface (Fig. 5b). The 2D nature of the central quadrupole field (Fig. 2(a,b)) is verified by the robust operation of the highly elongated 2D MOT (Fig. 3e). Freed from the slow field-switching constraints, we perform atomic spectroscopy at a repetition rate as high as $f_{\rm rep}=250$~kHz, utilizing nanosecond probe pulses as shown in Fig. 4. We note that the quasi-continuous spectroscopic measurements minimally perturb the atomic trajectories [17, 8]. In the ONF near field, the atomic density distribution is preserved as that during MOT operation, likely to be quite uniform. Therefore, our ONF spectroscopy should reflect the near-field shifts in a straightforward manner. As in Sec. III.5 and detailed in Appendix C, that the absorption profile is uniformly broadened without notable asymmetry [49] is quite unexpected and warrants further investigation.
The $\mu$-metal soft ferromagnetics are excellent field uniformizers, and can be precisely machined and positioned to generate exceptionally accurate 2D fields. In Sec. IV we provide arguments that a field-free ONF interface with $B_{1}\sim 100$ G/cm and $l\approx 100$~$m$ can be constructed using an $n=4$ assembly, where an ${\rm OD}\sim 10^{3}$ may be reached for an ONF-lattice [42] with 2D-MOT loading. The quantum optical measurements can be performed in a quasi-continuous manner, as those demonstrated in this work. The efficient laser-cooling in between measurements may help to mitigate ONF vibration induced heating [53]. The unprecedentedly high light-atom coupling strength may help to unlock new research opportunities in waveguide QED [45, 2, 4]. Bringing the ferromagnetic plates closer to the center, even higher gradients at $B_{1}\sim$ 1 kG/cm level appears achievable, where the ONF may serve as a local probe to study cooling-trapping dynamics in a high-gradient 2D-MOT [61]. Additional cooling in the high-gradient magnetic trap should lead to tightly confined 1D atomic samples, with which a natural step forward is to combine the ONF interface with magnetically guided atom interferometry [19] to explore controllable quantum interaction between the waveguided photons and atoms.
Acknowledgments
@@
DATA AVAILABILITY
Data and simulation codes underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
@@ Appendix A Magnetic field simulation
The accuracy of the field simulation relies on the granularity of magnet segmentation and the numerical relaxation tolerance. We adjusted these parameters to ensure that neither finer subdivisions nor tighter relaxation tolerances significantly affected the final results. The relaxation tolerance is set to ${\rm prec}=10^{-4}$ T (1 G). For the simulations in Fig. 2 and Fig. 5c, each NdFeB block is divided into $10\times 10\times 10$ subunits. Each $\rm{\bf\mu}$}-metal plate is divided into $n_{z}=160$ and $n_{y}=50$ segments along the $z$ (“$L$ direction”) and $y$ (“$W$ direction”) axes, respectively. Additional simulations confirmed that the central field profile $B_{y}(z)$, as shown in Fig. 5c, is insensitive to segmentation along the $x$-axis, allowing us to set $n_{x}=1$ to conserve computational resources. This approach neglects the variation of $B_{y}$ across the plate thickness ($x$), which is minimized at the center ($x=0$) and peaks at the edges. To account for this approximation, an overall scaling factor of 0.8 is applied to the field simulation in Fig. 5c, evaluated at $h=0.5~\mathrm{mm}$ from $S_{2}$, to match experimental measurements shown in Fig. 5b. This correction is necessary only in the vicinity of $S_{2}$. For field distributions where $h\gg d$, the simulation remains accurate, and we fix the NdFeB block surface field at $B_{0}=0.9~\mathrm{T}$ for all simulations.